# Unified Batch Normalization: Identifying and Alleviating the Feature Condensation in Batch Normalization and a Unified Framework

### 27 Nov 2023 v1

@@ 4.1 Implementation Details
On the ImageNet dataset. Consistent with common practices, we applied data augmentation and randomly cropped images to $224\times 224$ pixels for training ResNets [13], ResNeXt-50 [34], and Res2Net-50 [10]. ResNet-50 was trained using four Nvidia A100 GPUs and the other networks were trained using eight Nvidia A100 GPUs. Each model was trained for 90 epochs. We adopted a multi-step decay process for the learning rate. The optimizer was SGD, with an initial learning rate of 0.1, a momentum of 0.9, and a weight decay of 1e-4. The learning rate was programmed to decrease by 0.1 at the 30th and 60th epochs. The batch size per GPU was set to 256. In each model, the BN layer was substituted with our UBN, detailed in Table 2.
On the CIFAR-10 dataset. Our experiments were carried out on the CIFAR-10 dataset using three Nvidia 3090 GPUs. We adhered to standard data augmentation techniques, including random cropping of images to $32$ pixels for training ResNets [13] and VGGs [28]. Each model was trained for a total of 200 epochs. For learning rate management, we applied a multi-step decay strategy. Our optimization choice was SGD, starting with a learning rate of 0.1, a momentum of 0.9, and a weight decay rate of 1e-4. We programmed the learning rate to reduce by 0.1 at the 100th and 150th epochs. Each GPU was assigned a batch size of 128. In each model, the BN layer was replaced with our proposed UBN, as detailed in Table 3.
Figure 4: The learning curves and feature condensation curves of UBN and other normalization methods with ResNet-50 on the CIFAR-100 dataset. UBN can significantly reduce the feature condensation phenomenon without abandoning the batch statistics like GN [33], LN [1], and IN [31]. The leveraging of feature condensation threshold boost the stability of training, and the unification of rectifications of each component in BN boost the performance of UBN compared with existing normalization methods. We used non-uniform axis scale to visualize the differences between UBN and other state-of-the-art (SOTA) normalization methods in terms of loss and accuracy.
On the CIFAR-100 dataset. We also conducted experiments on the CIFAR-100 dataset using three Nvidia 3090 GPUs. Our process followed typical data augmentation practices, including random cropping of images to $32$ pixels for training ResNets [13], ResNeXts [34], VGGs [28], and Inception-v4s [29]. Each model underwent 200 training epochs. We introduced a warm-up phase in the first epoch, where the learning rate increased linearly with each batch to stabilize early training. After this, we employed a multi-step decay for the learning rate. Using SGD as the optimizer, we started with an initial learning rate of 0.1, a momentum of 0.9, and a weight decay of 5e-4. The learning rate was set to decrease by 0.2 at the 60th, 120th, and 160th epochs. We maintained a batch size of 128 per GPU. The BN layer in each model was replaced with our UBN, as shown in Table 4.
@@ 4.2 Performance Evaluation
Comparison of other normalization methods. On the CIFAR-100 dataset, we compared UBN with existing BN variants, as detailed in Table 4 and Figure 4. The batch size was set to 128 for each GPU. GN [33], which doesn’t utilize batch statistics, underperforms compared to standard BN in typical training setups. UBN, however, preserves the advantages of batch dependency while incorporating conditional adaptive running statistics. It also outperforms other leading normalization methods like RBN [11], IEBN [20], and BNET [35]. These methods tend to focus on specific aspects of BN, making targeted adjustments but overlooking a comprehensive perspective of BN from a running statistics viewpoint. UBN consistently demonstrates higher accuracy early in training, maintaining this advantage to the end.
Figure 6: Ablation of testing accuracy when applying UBN with different feature condensation threshold $\tau$. (a) The average feature cosine similarity curves of the features in the last batch of each epoch before the first normalization layer (b) The training loss curves (c) The testing accuracy curves. Using the threshold can alleviate the feature condensation problem. We show the average cosine similarity for a batch of samples at the end of each epoch for evaluation. We used non-uniform axis scale to visualize the differences between UBN and other state-of-the-art (SOTA) normalization methods in terms of loss and accuracy.
Incorporating into different network architectures. As a modification of the original Batch Normalization, our developed Unified Batch Normalization (UBN) is designed to replace BN layers in various networks, thereby improving classification performance across multiple datasets including ImageNet, CIFAR-100, and CIFAR-10, as indicated in Tables 2, 3, and 4. DNNs using UBN, particularly ResNets, demonstrate superior top-1 accuracies over those employing standard BN. On the ImageNet dataset, notable enhancements are observed in accuracies with ResNet-50 gaining a 3.27% improvement, ResNet-101 showing a 4.32% increase, and ResNext-50 demonstrating a 2.86% improvement. Res2Net-50 also achieving a 3.28% performance gain. On CIFAR-10 and CIFAR-100 datasets, UBN based ResNets exhibit the most notable improvement over BN based ResNets, achieving performance gains of 2.1% and 1.4% respectively. For VGG architectures, the implementation of UBN results in a 0.3% accuracy increase on both datasets, compared to BN based VGGs.
@@ 4.3 Ablation Study
Effects of each rectification in UBN. Our ablation studies focus on each rectification in UBN – normalization (including centering and scaling), affine, and running statistics rectification with feature condensation threshold. We conducted experiments on ResNet-50 and ResNet-101 on the CIFAR-10 dataset, as shown in Figure 5. For the experiment on ResNet-50, BN+Stats Rectification used threshold $\tau=0.25$ and UBN used threshold $\tau=0.15$. For experiments on ResNet-101, both BN+Stats Rectification and UBN used threshold $\tau=0.25$. We find that each rectification type improves BN performance. When combined, these modifications result in a surprisingly efficient performance gain throughout training. Using UBN also leads to higher accuracies after just one epoch of training.
Different feature condensation threshold $\tau$. We analyze the effect of feature condensation threshold $\tau$, which determines the running statistics in UBN. We show the results of the ResNet-101 model on the CIFAR-10 dataset. The specific statistics of each batch of samples is unpredictable, thus we can only provide a constant threshold to alleviate feature condensation. As shown in Figure 6, UBN with different threshold $\tau$ both boost the testing performance in the learning of DNNs. We demonstrate that the thresholds can dramatically alleviate the feature condensation which causes training efficiency in the early epochs. Although there is no linear or trivial relationships between the thresholds and the testing accuracies, the feature condensation can be alleviated by leveraging such threshold.
Figure 7: Effectiveness ablation (final testing accuracies) on ResNet-18, ResNet-34, and ResNet-50 trained on the CIFAR-10 dataset with different batch sizes for 200 epochs.
Different batch sizes. We examine UBN’s performance across different architectures on CIFAR-10, using ResNet-18, ResNet-34, and ResNet-50, as shown in Figure 7. We evaluated batch sizes of 32, 64, 128, 256, and 512. We observe that UBN’s testing accuracy decreases with smaller batch sizes, but the decline gets slower in deeper models.
@@ 5 Conclusion
Previous studies have mainly focused on improving certain Batch Normalization (BN) aspects in modern deep neural networks. In contrast, our work establish a unified approach to address this issue from a feature condensation perspective. Specifically, we present Unified Batch Normalization (UBN) framework, which involves a simple yet effective feature condensation rectification and precise adjustments to each component of BN. We rectify the feature condensation with a simple pre-defined threshold to boost the robustness of running statistics, and employ rectifications on each component of BN to improve the feature representation. Furthermore, UBN is designed to be easily integrated into existing models. By ensuring appropriate running statistics and applying comprehensive rectifications across all normalization components, UBN can significantly improve both testing performance and training efficiency.
Supplementary Material
Appendix A Exploring Feature Condensation in BN in Additional Experiments
@@
Appendix B Exploring Various Feature Condensation Threshold $\tau$ for UBN in Additional Experiments
@@
Appendix C Exploring the Effect of Various Batch Sizes for UBN in Additional Experiments
In this section, we conducted more experiments on different datasets and different models to compare the effect of different batch sizes in the training of UBN compared with traditional BN, which is shown in Figure 7 before. We conducted experiments on the ResNet-50 and ResNet-101, with batch size $32,64,128,256,512$, respectively. Extensive results of ResNet-50 on CIFAR-10 dataset are shown in Figure 18, Figure 19, Figure 20, Figure 21, and Figure 22. Extensive results of ResNet-101 on CIFAR-10 dataset are shown in Figure 23, Figure 24, Figure 25, Figure 26, and Figure 27. Extensive results of ResNet-50 on CIFAR-100 dataset are shown in Figure 28, Figure 29, Figure 30, Figure 31, and Figure 32. Extensive results of ResNet-101 on CIFAR-100 dataset are shown in Figure 33, Figure 34, Figure 35, Figure 36, and Figure 37.
@@ C.2 On the CIFAR-100 dataset
Appendix D Exploring the Functionality of Rectifications for UBN in Additional Experiments

### 08 Mar 2024 v2

@@ 4.1.2 On CIFAR-100
4.1.3 On CIFAR-10
Our experiments were carried out on the CIFAR-10 dataset using three Nvidia 3090 GPUs. We adhered to standard data augmentation techniques, including random cropping of images to $32$ pixels for training ResNets [13] and VGGs [30]. Each model was trained for a total of 200 epochs. For learning rate management, we applied a multi-step decay strategy. Our optimization choice was SGD, starting with a learning rate of 0.1, a momentum of 0.9, and a weight decay rate of 1e-4. We programmed the learning rate to reduce by 0.1 at the 100th and 150th epochs. Each GPU was assigned a batch size of 128. In each model, the BN layer was replaced with our proposed UBN, as detailed in Table 4.
Figure 4: The learning curves and feature condensation curves of UBN and other normalization methods with ResNet-50 on the CIFAR-100 dataset. UBN can significantly reduce the feature condensation phenomenon without abandoning the batch statistics like GN [35], LN [1], and IN [33]. The leveraging of feature condensation threshold boost the stability of training, and the unification of rectifications of each component in BN boost the performance of UBN compared with existing normalization methods. We used non-uniform axis scale to visualize the differences between UBN and other SOTA normalization methods in both loss and accuracy.
Table 4: Classification performance of different normalization methods on the CIFAR-10.
CIFAR-10 Normalization Top-1 Accuracy (%) Top-5 Accuracy (%)
ResNet-18 BN 93.65 99.85
UBN($\tau=0.025$) 93.80 99.81
ResNet-34 BN 93.02 99.76
UBN($\tau=0.15$) 93.98 99.83
ResNet-50 BN 92.26 99.84
UBN($\tau=0.15$) 94.32 99.85
ResNet-101 BN 93.19 99.87
UBN($\tau=0.25$) 94.56 99.89
VGG-11 BN 91.14 99.65
UBN($\tau=0.25$) 91.48 99.66
@@ 4.2 Performance Evaluation
4.2.1 Comparison of other normalization methods
On the CIFAR-100 dataset, we compared UBN with existing BN variants, as detailed in Table 3 and Figure 4. The batch size was set to 128 for each GPU. GN [35], which doesn’t utilize batch statistics, underperforms compared to standard BN in typical training setups. UBN, however, preserves the advantages of batch dependency while incorporating conditional adaptive running statistics. It also outperforms other leading normalization methods like RBN [11], IEBN [20], and BNET [37]. These methods tend to focus on specific aspects of BN, making targeted adjustments but overlooking a comprehensive perspective of BN from a running statistics viewpoint. UBN consistently demonstrates higher accuracy early in training, maintaining this advantage to the end.
4.2.2 Incorporating into different network architectures
As a modification of the original Batch Normalization, our developed Unified Batch Normalization (UBN) is designed to replace BN layers in various networks, thereby improving classification performance across multiple datasets including ImageNet, CIFAR-100, and CIFAR-10, as indicated in Tables 2, 4, and 3. DNNs using UBN, particularly ResNets, show superior top-1 accuracies over those employing standard BN. On ImageNet, notable enhancements are observed in accuracies with ResNet-50 gaining a 3.27% improvement, ResNet-101 showing a 4.32% increase, and ResNext-50 with a 2.86% improvement. Res2Net-50 also achieving a 3.28% gain. On CIFAR-10 and CIFAR-100, UBN based ResNets exhibit the most notable improvement over BN based ResNets, achieving performance gains of 2.1% and 1.4% respectively. For VGG architectures, the implementation of UBN results in a 0.3% accuracy increase on both datasets, compared to BN based VGGs.
4.2.3 Generalization to Typical Vision Tasks
Our proposed UBN method is designed to enhance the representational capabilities of models across a variety of tasks, effectively serving as a superior alternative to traditional BN. We evaluated UBN’s performance on typical vision tasks by integrating it into Faster-RCNN for object detection and Mask-RCNN for instance segmentation within the COCO dataset. Each model underwent training for a duration of 12 epochs, utilizing a multi-step decay strategy for optimal learning rate adjustment. The optimization was carried out using SGD with an initial learning rate set at 0.02, momentum at 0.9, and weight decay at 1e-4. We used a batch size of 2 for each GPU (8 in total) during the training process. The empirical results, as detailed in Table 5 and Table 6, demonstrate that UBN substantially surpasses BN, achieving an improvement exceeding 4% in mean Average Precision (AP).
Table 5: Object Detection Results on COCO using Faster-RCNN
[…]
ResNet-50-BN 33.1 53.2 35.6 18.4 35.8 43.6
ResNet-50-UBN ($\tau=0.15$) 37.7 58.6 40.5 20.8 40.9 49.5
ResNet-101-BN 31.9 51.4 34.1 17.6 34.5 42.2
ResNet-101-UBN ($\tau=0.15$) 37.6 58.2 40.1 20.7 40.7 50.4
Table 6: Instance Segmentation Results on COCO using Mask-RCNN
[…]
ResNet-50-BN 34.0 54.0 36.6 18.5 36.9 44.5
ResNet-50-UBN ($\tau=0.15$) 38.5 59.3 41.6 21.3 41.9 50.5
ResNet-101-BN 33.1 52.5 35.7 17.8 35.8 44.0
ResNet-101-UBN ($\tau=0.15$) 38.6 59.1 41.4 22.3 41.6 51.8
[…]
ResNet-50-BN 31.1 50.8 33.0 13.5 33.1 46.4
ResNet-50-UBN ($\tau=0.15$) 34.8 56.1 36.8 15.6 37.3 50.7
ResNet-101-BN 30.3 49.4 32.0 12.7 31.8 45.7
ResNet-101-UBN ($\tau=0.15$) 34.6 55.8 36.6 16.5 36.6 51.2
@@ 4.3 Ablation Study
4.3.1 Effects of each rectification in UBN
Our ablation studies focus on each rectification in UBN – normalization (including centering and scaling), affine, and running statistics rectification with feature condensation threshold. We conducted experiments on ResNet-50 and ResNet-101 on the CIFAR-10 dataset, as shown in Figure 5. For the experiment on ResNet-50, BN+Stats Rectification used threshold $\tau=0.25$ and UBN used threshold $\tau=0.15$. For experiments on ResNet-101, both BN+Stats Rectification and UBN used threshold $\tau=0.25$. We find that each rectification type improves BN performance. When combined, these modifications result in a surprisingly efficient performance gain throughout training.
4.3.2 Different feature condensation threshold $\tau$
We analyze the effect of feature condensation threshold $\tau$, which determines the running statistics in UBN. We show the results of the ResNet-101 model on the CIFAR-10 dataset. The specific statistics of each batch of samples is unpredictable, thus we can only provide a constant threshold to alleviate feature condensation. As shown in Figure 6, UBN with different threshold $\tau$ both boost the testing performance in the learning of DNNs. We demonstrate that the thresholds can dramatically alleviate the feature condensation which causes training efficiency in the early epochs. Although there are no linear or trivial relationships between the thresholds and the testing accuracies, feature condensation can be alleviated by leveraging such a threshold. For simplicity and efficiency in our experimental setup, we empirically set the threshold $\tau=0.15$. This choice, made without manual tuning, worked effectively across the majority of our tests.
Figure 6: Ablation of testing accuracy when applying UBN with different feature condensation threshold $\tau$. (a) Average feature cosine similarity curves of the features in the last batch of each epoch before the first normalization layer (b) Training loss curves (c) Testing accuracy curves. Using the threshold can alleviate the feature condensation problem. We show the average cosine similarity for a batch of samples at the end of each epoch. We used non-uniform axis scale to visualize the differences between UBN and other normalization methods w.r.t. loss and accuracy.
4.3.3 Different batch sizes
We examine UBN’s performance across different architectures on CIFAR-10 and CIFAR-100. We conducted experiments on ResNet-18, ResNet-34, and ResNet-50, as shown in Table 7. We evaluated batch sizes of 32, 64, 128, 256, and 512. We did not observe distinct correlation between batch size and test accuracy.
Table 7: Effectiveness ablation (final testing accuracies) on ResNet models trained on CIFAR-10 and CIFAR-100 datasets with different batch sizes for 200 epochs.
Dataset Model Batch Size
32 64 128 256 512
CIFAR-10 ResNet-18 94.45 94.90 94.32 93.79 93.38
ResNet-34 95.51 94.73 94.54 94.34 93.78
ResNet-50 95.39 94.59 94.58 94.57 94.13
CIFAR-100 ResNet-50 78.02 78.52 78.99 78.86 77.21
ResNet-101 77.53 78.27 79.33 78.03 77.87
@@ 5 Conclusion
Previous studies have mainly focused on improving certain Batch Normalization (BN) aspects. In contrast, our work establishes a unified approach from a feature condensation perspective. Specifically, we present Unified Batch Normalization (UBN) framework, which involves a simple yet effective feature condensation rectification and precise adjustments to each component of BN. We rectify the feature condensation with a simple pre-defined threshold to boost the robustness of running statistics, and employ rectifications on each component of BN to improve the feature representation. Furthermore, UBN can be easily integrated into existing models. By ensuring appropriate running statistics and applying comprehensive rectifications across all normalization components, UBN can significantly improve both testing performance and training efficiency.
Limitations: Feature condensation occurs in different scenarios throughout the training of DNNs, making the precise determination of a feature condensation threshold challenging. While UBN significantly reduces training duration, the necessity for manual adjustment of this threshold and the computation of feature similarity may cause additional time and computational cost.
Appendix 0.A Exploring Feature Condensation in BN in Additional Experiments
@@
Appendix 0.B Exploring Various Feature Condensation Threshold $\tau$ for UBN in Additional Experiments
@@
Appendix 0.C Exploring the Effect of Various Batch Sizes for UBN in Additional Experiments
In this section, we conducted more experiments on different datasets and different models to compare the effect of different batch sizes in the training of UBN compared with traditional BN, which is shown in Figure LABEL:fig:ablation-bs before. We conducted experiments on the ResNet-50 and ResNet-101, with batch size $32,64,128,256,512$, respectively. Extensive results of ResNet-50 on CIFAR-10 dataset are shown in Figure 17, Figure 18, Figure 19, Figure 20, and Figure 21. Extensive results of ResNet-101 on CIFAR-10 dataset are shown in Figure 22, Figure 23, Figure 24, Figure 25, and Figure 26. Extensive results of ResNet-50 on CIFAR-100 dataset are shown in Figure 27, Figure 28, Figure 29, Figure 30, and Figure 31. Extensive results of ResNet-101 on CIFAR-100 dataset are shown in Figure 32, Figure 33, Figure 34, Figure 35, and Figure 36.
@@ 0.C.2 On the CIFAR-100 dataset
Appendix 0.D Exploring the Functionality of Rectifications for UBN in Additional Experiments
@@
Appendix 0.E Analysis of training complexity
We’ve conducted new experiments to evaluate UBN’s training efficiency by calculating the comparative training time when acheving the comparative training performance for each model. The reported UBN based model results achieve similar performance levels to BN based model. Although determination of a feature condensation threshold may cause additional computational cost and time, UBN reaches comparable performance in less time, as shown in Table 8.
Table 8: Training time of BN and UBN on different models on CIFAR-10 and ImageNet datasets.
Dataset Model Training Time (s) #Epoch
CIFAR-10 ResNet-34-BN 3760 200
ResNet-34-UBN ($\tau=0.15$) 3605 101
ResNet-50-BN 4302 200
ResNet-50-UBN ($\tau=0.15$) 3606 73
VGG-11-BN 1475 200
VGG-11-UBN ($\tau=0.25$) 1044 114
ImageNet Res2Net-50-BN 33048 90
Res2Net-50-UBN ($\tau=0.15$) 23734 31
ResNeXt-50-BN 29367 90
ResNeXt-50-UBN ($\tau=0.15$) 22840 31
ResNet-101-BN 35594 90
ResNet-101-UBN ($\tau=0.15$) 26557 31